kZAvala:

Angle RST is a right angle. Angle RSU has a measure of 25°. What is the measure of angle TSU? 25° 45° 65° 75°

Shadow:

hello @kZAvala

kZAvala:

hello

Shadow:

Is there an image of the shape associated with this problem?

fall15:

I think he means that ^

kZAvala:

1 attachment
Shadow:

There we go

kZAvala:

can you see it?

Shadow:

Yes

Shadow:

RST = 90 RSU = 25 TSU = RST - RSU TSU = 90 - 25 TSU = 65

kZAvala:

thank you !

fall15:

That was quick. . .

Shadow:

RSU is basically a part of the angle RST. When you subtract angle RSU from angle RST, you are left with the measure for the angle TSU

kZAvala:

oooh okay